OpenStudy (anonymous):

The ratio of the angle measures in a quadrilateral is 1 : 4 : 5 : 6. Find the measure of the largest angle. A. 135° B. 145° C. 90° D. 112.5°

OpenStudy (anonymous):

let x be the ratio 1x + 4x + 5x + 6x = 360 (because the quadrilateral's interior angles sums up to 360) 16x = 360 divide both sides by 16 to get x x = 22.5 Substituting it to the biggest number in the ratio, we have 6x = 6(22.5) = 135 therefore, the answer is A.

OpenStudy (anonymous):

1x+ 4x+ 5x+ 6x=360 16x=360 x=360/16 x=22.5 6*22.5=135 is : A. 135°